OpenStudy (anonymous):

2x+1 over 3=13 over 3 wats x

OpenStudy (anonymous):

what does the equation say?

OpenStudy (anonymous):

(2x + 1)/3 = 13/3 2x +1 = 13 * 3 / 3 2x + 1 = 13 (as 3 and 3 get cancelled ) 2x = 13 - 1 2x = 12 x = 12/2 x = 6